OpenStudy (anonymous):

Evaluate the function at each specified value of the independent variable and simplify. (If an answer is undefined, enter UNDEFINED.) h(t) = t2 − 2t

OpenStudy (turingtest):

specify some values...

myininaya (myininaya):

for example if we have t=0, we have wherever there is a t, we plug in zero \[h(0)=0^2-2(0)=0-0=0\] another example if we have t=1, we have wherever there is a t, we plug in one \[h(1)=1^2-2(1)=1-2=-1\] last example if we have t=-3 we have wherever there is a t, we plug in negative three \[h(-3)=(-3)^2-2(-3)=9-(-6)=9+6=15\] -- Do you think polynomials are undefined anywhere?